Radison Enterprises sells a product for $98 per unit. The variable cost is $50 per unit, while fixed costs are $387,072.

Determine (a) the break-even point in sales units and (b) the break-even point if the selling price were increased to $106 per unit.

a. Break-even point in sales units units
b. Break-even point if the selling price were increased to $106 per unit units
Answer #1

>> Break even point in sales units = ( Fixed cost ) / Contribution margin per unit.

>> Contribution margin per unit = Selling price - Variable cost per unit.

>> Contribution margin per unit = $ 98 - $ 50 = $ 48.

>> Break even point in sales units = $ 387,072 / $ 48 = 8,064.

___________________________________________________________________________________________

>> New Contribution margin per unit = $ 106 - $ 50 = $ 56.

>> New Break even point in sales units = $ 387,072 / $ 56 = 6,912.

____________________________________________________________________________________________
Answer #2

SOLUTION :


a.



Contribution margin per unit (UCM) = SP - UVC = 98 - 50 = 48 ($)


BEP (units) = FC / UCM = 387072 / 48 = 8064 units (ANSWER)





b.


If SP = $106 :


Contribution margin per unit (UCM) = SP - UVC = 106 - 50 = 56 ($)


BEP (units) = FC / UCM = 387072 / 56 = 6912 units (ANSWER)

answered by: Tulsiram Garg